OpenStudy (anonymous):

Both of these questions have to deal with using inverses to linear systems: 1. Work Schedule You worked 18 hours last week and earned a total of $124 before taxes. Your job as a lifeguard pays $8 per hour, and your job as a cashier pays $6 per hour. How many hours did you work at each job? 2. Law Enforcement During one calendar year, a state trooper issed a total of 385 citations for warning and speeding tickets. Of these, there were 31 more warnings than speeding tickets. How many warnings and how many speeding tickets were issued?

OpenStudy (callisto):

For the first question, I think you need to set up the equation first. Let x be the number of hours you worked as a lifeguard, then the number of hours you worked as a cashier is ...??

hero (hero):

Hint: (For the first one) x = life guard hours y = cashier hours x + y = 18 8x + 6y = 124(18) I don't get what they mean by "inverses" to linear systems

hero (hero):

@Callisto, do you have a clue what they mean by that?

OpenStudy (callisto):

Actually, I have no ideas about that. But we can solve the problems by setting up equation. Btw, since the total working hour is 18, when we let x be the working hours as the lifeguard, working hours as a cashier can be found be subtracting x from 18. In this way, we just need one equation.

hero (hero):

I already know that. I just prefer setting up a system of equations. The reason is so that if I ever need to, I can always just graph both equations and check the values that way.

hero (hero):

I just wanted to know what they meant by "inverse" linear system

OpenStudy (callisto):

Got it!